OpenStudy (veeveeniko):

What are the masses of oxygen and magnesium in 5.0 grams of MgO?

OpenStudy (veeveeniko):

@wolf1728 @butterflydreamer @IrishBoy123 Could anyone help me?

OpenStudy (veeveeniko):

@ganeshie8 @dan815 @Luigi0210 please help?

OpenStudy (veeveeniko):

@ikram002p @AaronAndyson @Ashley10116 @Bananas1234 @death_lover @Dbzfan836 @georgia545 @hunter2kb @hartnn @Jack1 @mattyboyy @omarbirjas @potatoprincess22 @rosedewittbukater @School5678 @SolidIceCream @TheKylieeM @xo_kansasprincess_xo @youngamerican Need all the help I can get

OpenStudy (mattyboyy):

I can try to help if you give me one moment. I'm doing economics work and I'm almost done.

OpenStudy (veeveeniko):

@mattyboyy Thank I'd appreciate it

OpenStudy (mattyboyy):

Okay I'm done

OpenStudy (veeveeniko):

How should I go about doing this problem

OpenStudy (mattyboyy):

4:30 in the morning and I'm trying to do chemistry -.- Honestly my mind is so off right now and I haven't done chemistry since last year..

OpenStudy (mattyboyy):

Have you looked on any websites for references?

OpenStudy (veeveeniko):

It's 4:30 for me too, trust me I know the feel.

OpenStudy (veeveeniko):

This is the only question I'm really stuck on

OpenStudy (mattyboyy):

http://faculty.uml.edu/james_hall/84123/5.htm Would this help?

OpenStudy (mattyboyy):

Or maybe @wolf1728 ?

OpenStudy (wolf1728):

Yeah I just got here

OpenStudy (mattyboyy):

http://www.ausetute.com.au/molreact.html Im honestly just tossing out websites right now cause im about to pass out lol

OpenStudy (wolf1728):

Atomic Mass of Mg 24.31 Oxygen 16

OpenStudy (veeveeniko):

How did you come to that conclusion @wolf1728 I want to be able to do the problem if it ever comes up again. Also @mattyboyy it's fine, thank you for at least attempting to help me. I appreciate it.

OpenStudy (wolf1728):

Mg is 1.519375 times more "massive" than oxygen

OpenStudy (wolf1728):

I was just listing their atomic masses

OpenStudy (veeveeniko):

So how do I solve the problem then?

OpenStudy (wolf1728):

MgO is one part Mg and one part O If they "weighed" the same then it would be 2.5 grams of each Since Mg "weighs" about 1.52 times more than O we need 1.52 more times of oxygen's mass Am I right?

OpenStudy (veeveeniko):

So you're saying Magnesium would be 4.02 and Oxygen would be 0.98?

OpenStudy (wolf1728):

No - I'm working on that right now.

OpenStudy (wolf1728):

Mg - 1.52 O - 1 Total = 2.52 Mass of Mg = 5* 1.52/2.52 = 3.0158730159 grams Mass of O = 5 * 1/2.52 = 1.9841269841 grams

OpenStudy (wolf1728):

I think that's right but it MIGHT be that Mg is 1.9841269841 grams and O is 3.0158730159 grams but I think my first answer is correct.

OpenStudy (veeveeniko):

Okay, I really appreciate your help.

OpenStudy (wolf1728):

Okay and thank you. :-)

OpenStudy (veeveeniko):

I understood how you came to that conclusion

OpenStudy (wolf1728):

Okay then